OpenStudy (anonymous):

Rewrite in simplest radical form 1 x −3 6 . Show each step of your process.

OpenStudy (decentnabeel):

\[\frac{\frac{1}{x}}{6^{-3}}=\frac{216}{x}\]

OpenStudy (decentnabeel):

\[\mathrm{Apply\:exponent\:rule}:\quad \:a^{-b}=\frac{1}{a^b}\] \[6^{-3}=\frac{1}{6^3}\] \[=\frac{\frac{1}{x}}{\frac{1}{6^3}}\] \[\mathrm{Divide\:fractions}:\quad \frac{\frac{a}{b}}{\frac{c}{d}}=\frac{a\cdot \:d}{b\cdot \:c}\] \[=\frac{6^3}{x}\]

OpenStudy (decentnabeel):

are you understand @boots_2000

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

not at all

OpenStudy (anonymous):

1 over x raised 3 over 6 thats what the question was supposed to be

Nnesha (nnesha):

\[\huge\rm \frac{ 1 }{ x^\frac{ 3 }{ 6 } }\] like this ?

OpenStudy (anonymous):

yes

Nnesha (nnesha):

alright you can reduce the fraction 3/6 =?

OpenStudy (anonymous):

1/2

Nnesha (nnesha):

yep right now you can change 1/2 root to radical \[\huge\rm x^\frac{ m }{ n } = \sqrt[n]{x^m}\]

OpenStudy (anonymous):

so my answer would be square root x of 1?

Nnesha (nnesha):

nope

Nnesha (nnesha):

\[\huge\rm x^\frac{ 1 }{ 2}=???\] let m = 1 and n =2 look at the exapmle i gave you

Nnesha (nnesha):

example*

Nnesha (nnesha):

yesright

Nnesha (nnesha):

sorry i didn't see word square so now you are not allowed to have radical at the denominator

OpenStudy (decentnabeel):

\[\frac{1}{x^{\frac{3}{6}}}=\frac{1}{\sqrt{x}}\] \[\mathrm{Simplify}\:\frac{3}{6}:\quad \frac{1}{2}\] =1/x^(1/2) \[=\frac{1}{\sqrt{x}}\]

OpenStudy (anonymous):

\[\sqrt[2]{x ^{1}}\]

Nnesha (nnesha):

\[\huge\rm \frac{ 1 }{ \sqrt{x}}\] multiply both the denominator and numerator by square root of x

Nnesha (nnesha):

don't forget the 1 at the numerator that stay there

Nnesha (nnesha):

\[\huge\rm \frac{ 1 }{ \sqrt{x} } \times \frac{ \sqrt{x} }{ \sqrt{x} }\]

OpenStudy (anonymous):

\[1\div \sqrt[2]{x ^{1}}\]

OpenStudy (anonymous):

like that?

Nnesha (nnesha):

yes right now multiply both the top and bottom by the denominator (sqrt{{x}) = answer

OpenStudy (anonymous):

\[-\sqrt[2]{x ^{1}}\]

Nnesha (nnesha):

nope how did you get negative sign or is it typo ? ;)

OpenStudy (anonymous):

typo

Nnesha (nnesha):

btw radical sign mean square root so you don't have to write 2 .....

OpenStudy (anonymous):

sorry

Nnesha (nnesha):

\[\sqrt{ }\] <-- square root

Nnesha (nnesha):

\[\huge\rm \frac{ 1 }{ \sqrt{x} } \times \frac{ \sqrt{x} }{ \sqrt{x} }\] multiply denominator by denominator and numerator y numerator

OpenStudy (anonymous):

so the answer is without the negative sign???

OpenStudy (anonymous):

so its just the suare root of x that we just put but withour the negative???

Nnesha (nnesha):

well there isn't any negative sign in the original question so it's pretty obvious :=) o^_^o

Nnesha (nnesha):

nope multiply

Nnesha (nnesha):

\[\huge\rm \frac{ 1 }{ \sqrt{x} } \times \frac{ \sqrt{x} }{ \sqrt{x} }\] \[\frac{ 1 \times \sqrt{x} }{ \sqrt{x} \times \sqrt{x}}\]

OpenStudy (anonymous):

okay well im lost so imma just guess or not answer it

Nnesha (nnesha):

as you wish.. just one last step MULTIPLCATION! that's it done!